# Pretraining Frame Preservation in Autoregressive Video Memory Compression

### v1

@@ 1 Introduction
Storytelling capability, narrative coherence, and context quality are increasingly important in the video generation community. The latest commercial models like Sora2 [sora2], Veo3.1 [veo31], and Kling2.5 [kling25] enable storyboard creation-based, scene planning, and dynamic camera shots (e.g., in the style of TikTok or YouTube Shorts). Recent academic models have also focused on long video streaming [yin2024slow, huang2025selfforcing], scene planning [guo2025lct, huang2024context, zhao2024moviedreamer, zheng2024videogen, zhou2024storydiffusion], and video context consistency [huang2025conceptmaster, jiang2024videobooth, long2024videostudiogeneratingconsistentcontentmultiscene]. Among these approaches, autoregressive models are a key paradigm for video storytelling, supporting native video continuation and storyboard streaming. Autoregressive models view video history as context, and they face unique challenges when handling long-form content like in-the-wild videos and movies.
The trade-off between context quality and length is a typical and critical problem in video context modeling. A naive sliding window, which cuts off all distant-enough frames, maintains a fixed context length but loses long-range history. Various compression approaches have been extensively studied; e.g., highly compressed VAEs like LTXV [hacohen2024ltx] and DC-AE [chen2024deep] show that it is possible to compress pixel-space videos into more compact contexts, and hybrid approaches [bachmann2025flextok, jin2024pyramidal, zhou2025hitvideo] like FramePack [zhang2025framepack] show that compression can happen at multiple levels, albeit at the cost of high-frequency image details. Another strategy is to preserve the context length but reduce computation (e.g., using sparse [xi2025sparse, zhang2025spargeattn, zhang2025sta, xia2025trainingfree] or linear [cai2023efficientvit, xie2024sana, wang2020linformer, choromanski2020rethinking, yu2022metaformer, katharopoulos2020transformers] attention), though the linear layer cost remains an overhead for training and bidirectional inference. More specialized methods like token merging [bolya2023tomesd, bolya2023tome] also demonstrate that a higher merging rate leads to greater detail loss. These research efforts implicitly converge on the observation that context quality and length are governed by a fundamental trade-off where the optimal balance has not yet been reached.
@@
Experiments show that the pretrained memory encoder can, by itself, reconstruct history frames at high quality (measured by PSNR/SSIM [Wang2004Image]). When connected to video diffusion models for autoregressive finetuning, the encoder provides long-range consistency. This framework achieves a balanced and optimized point in the context length-quality trade-off, leading to a practical autoregressive video model with a relatively short context length and perceptually high consistency based on its long-range history.
@@ 2 Related Work
Autoregressive video diffusion and long videos. Diffusion has become the driving force behind video synthesis, where a central challenge is length scaling. Training-free length‑extension methods [qiu2023freenoise, lu2025freelong, ma2025freelongpp, zhao2025riflex, ruhe2024rolling] reschedule noise or re-balance temporal frequency to stretch pre-trained models beyond their training horizon. A complementary thread blends diffusion with causal prediction: Diffusion Forcing [chen2025diffusion] and HistoryGuidance [song2025history] enable variable-horizon conditioning and stable long rollouts by noise injection. These approaches are adapted in industrial systems such as SkyReels‑V2 [chen2025skyreelsv2infinitelengthfilmgenerative] and Magi-1 [teng2025magi1]. Additionally, StreamingT2V [henschel2024streamingt2v] augments existing models with short- and long-term memory with randomized blending to generate longer videos. FAR [gu2025longcontextautoregressivevideomodeling] studies long‑context AR with flexible RoPE [su2024roformer] decay and mixed short/long windows, while CausVid [yin2024slow] distills a bidirectional teacher to a few‑step causal generator. StreamDiT [kodaira2025streamdit] combines multi-step distillation with a moving frame buffer and mixed partition training to generate results in real-time. To mitigate error accumulation during AR generation, Self‑Forcing [huang2025selfforcing] simulates AR rollout during training, while its extensions [cui2025selfforcingpp, liu2025rollingforcing, yang2025longlive] further improve length generalization.
@@
Context learning and compression. Long-context persistence is crucial as we extend video generation beyond a few seconds. One option is retrieval, which grounds generation in a persistent state. WorldMem [xiao2025worldmem] and Context-as-Memory [yu2025contextasmemory] augment AR video world models with FoV-based history retrieval, while VMem [li2025vmem] indexes past views by surfels to retrieve the most relevant viewpoints. Memory Forcing [huang2025memoryforcing] pairs geometry-indexed spatial memory with tailored training regimes to balance exploration and revisits. Beyond fixed retrieval rules, Mixture-of-Contexts [cai2025moc] learns a dynamic sparse attention context router so that tokens attend only to the most salient chunks. Similarly, MoGA [jia2025moga] and Holocine [meng2025holocine] also propose token-level sparse attending policies. Pack-and-Force [wu2025packandforce] instead proposes a learnable context semantic retriever. Orthogonal to retrieval, compression of history turns an unbounded context into a compact state. FramePack [zhang2025framepack] compresses prior frames into a fixed‑size latent “packed” context. Captain Cinema [xiao2025captaincinema] uses a similar compression for keyframes. StateSpaceDiffuser [savov2025statespacediffuser] and Po et al. [po2025ssmworldmodel] swap quadratic attention for recurrent states to maintain long‑term memory. TTTVideo [dalal2025oneminutevideogenerationtesttime] and LaCT [zhang2025lact] use light MLP test-time training layers as a learned context representation.
@@
Efficient video diffusion designs. As we scale video generation to long horizons, large context windows are bottlenecked, driving a wave of efficient computational designs. Kernel advances such as FlashAttention [dao2022flashattentionfastmemoryefficientexact, dao2023flashattention2] improve throughput. Static or hardware-friendly sparse patterns include sliding/tiling 3D windows [zhang2025sta], radial spatiotemporal masks [li2025radialattention], and training-free head/pruning heuristics [xi2025sparse, yang2025svg2]. Dynamic or learned pruning/routing further select salient token pairs or blocks [wu2025vmoba], coarse-to-fine sparse token selection [zhang2025vsa], Sage/SpargeAttention families [zhang2025sageattention, zhang2024sageattention2, zhang2025spargeattn], blockified routing with cached search [xia2025trainingfree], and progressive block carving [zhang2025jenga]. Another line of work leverages compressing the latent space or token sequence: token merging and patch scaling [bolya2022tome, lee2024tokenmergevid], compact/variable-rate tokenizers [bachmann2025flextok], highly compressed latent space [hacohen2024ltx], or multiscale pyramids with re-noising [jin2024pyramidal]. SANA [xie2024sana] introduced a linear‑attention Diffusion Transformer for images, while SANA‑Video [chen2025sanavideo] extends this with block‑linear attention and a constant‑memory KV cache.
@@ 3 Method
Preliminaries. We clarify the notation of a typical autoregressive video diffusion model. Unless otherwise noted, all “frames”, “pixels”, etc., refer to latent concepts. We consider the typical Diffusion Transformers (DiTs) like Wan [wang2025wan] and HunyuanVideo [kong2024hunyuanvideo] with flow match scheduling. Rectified-flow models map noisy latents $\bm{X}_{t_{i}}\in\mathbb{R}^{T\times H\times W\times C}$ with
@@
where $\bm{c}$ is a set of conditions such as text prompts (a single prompt or streamed from a storyboard), and $t_{i}\sim\mathcal{L}(0,1)$ is the shifted logit-normal distribution [esser2024scaling] for flow matching.
@@ 3.2 Finetuning Video Diffusion Model
Figure 5: Visual comparison on compression reconstruction. We present reconstruction results after the pretraining using different possible neural structures and various compression settings. * The “Large Patchifier” is technically equivalent to [zhang2025framepack].
@@ 4.1 Experimental Details
Base models. All experiments are conducted on base models of HunyuanVideo [kong2024hunyuanvideo] and Wan [wang2025wan] family. For Wan 2.2, the involved models are the 5B model and the 14B high noise model. Only the high noise model is used to simplify experiments (similar to the design choice of WanAnimate).
Data preparation. The dataset is prepared with about 5 million internet videos from a range of websites. About half of the data is vertical Short-style videos and the rest are common horizontal videos. The data is cleaned for quality, and then the high-quality part is captioned with Gemini-2.5-flash VLM and the remainder is processed with local VLMs like QwenVL [wang2024qwen2]. The generated captions are in storyboard format with timestamps. When training the autoregressive models, only the prompt with the nearest timestamp is received by the model.
Storyboards. If not noticed, all storyboard prompts in this paper are written by Gemini-2.5-pro. We also tested open source alternatives like Qwen [wang2024qwen2] and the quality difference is negligible. See also the supplement for instructions.
@@ 4.2 Ablation Study
Table 1: Quantitative results on compression structures. We show numerical tests with different ablative compression architectures. * The “Large Patchifier” is technically equivalent to [zhang2025framepack].
@@
Compression model architecture. As shown in Fig. 5, we compare the image reconstruction details of different architectures under various compression rates, focusing on the performance of Only LR, Without LR, Large Patchifier, and our proposed method. The Only LR variant keeps only the low-resolution branch in Fig. 3, while Without LR removes it, leaving only the high-resolution branch. The Large Patchifier method increases the compression rate by enlarging the DiT’s patchifying projection kernel without introducing new layers, equivalent to FramePack [zhang2025framepack] and we directly use its implementation. The results indicate that the Only LR and Without LR variants produce reconstructions that differ significantly from the original images. The Large Patchifier method yields substantial structural changes. The proposed method, even at a high compression rate such as $4\times 4\times 2$, can still preserve much of the image appearance. A lower compression rate, such as $2\times 2\times 1$, provides the best detail preservation, albeit at the cost of a longer context length.
@@
Error accumulation (drifting). The problem of error accumulation highly depends on the training dataset. Surprisingly, we find that when trained on videos with dense camera shot shifts, like Short-style videos, the error accumulation problem seems to be absent or invisible. For other types of videos requiring long continuation over single shots, multiple design choices [zhang2025framepack, huang2025selfforcing, chen2025diffusion] are discussed in the supplement.
@@ 4.4 Quantitative Results
Metrics. We include several metrics for video evaluations from VBench [huang2023vbench], VBench2 [zheng2025vbench2], etc., with some modifications. Cloth: The cloth consistency by asking VLM questions. We upgrade VBench2 [zheng2025vbench2]’s LLaVA to Gemini-2.5-pro; Instance: The object-level quality by asking VLM questions about unnatural or inconsistent instances. We upgrade VBench2 [zheng2025vbench2]’s Qwen to Gemini-2.5-pro; Identity: The facial similarity using ArcFace [deng2019arcface] with face detection RetinaFace [deng2020retinaface]; Clarity: The MUSIQ [ke2021musiq] image quality predictor trained on SPAQ [fang2020perceptual]. This metric evaluates artifacts like noises; Aesthetic: The LAION aesthetic predictor [schuhmann2022laion], measuring the aesthetic values perceived by a CLIP estimator; Dynamic: The RAFT [teed2020raft] modified by VBench to estimate the degree of dynamics. Note that multi-shot videos intuitively have a special RAFT flow between shots, but in practice, RAFT seems to output near-zero values between shots, making the difference negligible from regular videos; Semantic: The video-text score computed by ViCLIP [wang2023internvid]. This metric measures the overall semantic consistency and prompt alignments.
@@
WanI2V [wang2025wan] + QwenEdit [wu2025qwenimagetechnicalreport] (1p) 94.10 68.45 85.21 /
WanI2V [wang2025wan] + QwenEdit [wu2025qwenimagetechnicalreport] (2p) 95.09 68.22 91.19 1198
WanI2V [wang2025wan] + QwenEdit [wu2025qwenimagetechnicalreport] (3p) 94.28 67.30 80.34 /
@@
Consistency analysis. We compare various architectures of our method, including different compression rates and ablations with or without pre-training and the LR branch. We also compare against a common baseline that combines Wan-2.2-I2V [wang2025wan] with Qwen-Image-Edit-2509 [wu2025qwenimagetechnicalreport]. This baseline first generates a series of initial frames with an image model, animates them into video clips, and then concatenates the clips. We test this baseline with 1, 2, and 3 images for Qwen-Edit as history contexts (since the model supports at most 3 inputs). Our evaluation focuses on the consistency of clothing, identity, and objects. For several practical and well-performed models, we also conducted a user study and evaluated their ELO scores.
@@ 4.5 Alternative Architecture
HunyuanVideo [kong2024hunyuanvideo] 12.8B ($\times4$$\times$2) 61.27 67.49 71.22 26.29 1189
Wan [wang2025wan] 2.2 14B* ($\times4$$\times$2) 67.22 69.37 69.81 27.12 1231
Wan [wang2025wan] 2.2 5B ($\times4$$\times$2) 66.25 69.01 65.13 25.99 1215
Wan [wang2025wan] 2.2 5B ($\times2$$\times$2) 66.37 68.95 66.29 26.13 1224
@@
Cross attention enhancement. As shown in Fig. 9, we demonstrate that connecting the features from the penultimate layer of the encoder to each block of the DiT via cross-attention (like IP-Adapter [ye2023ip]) can further enhance the model’s consistency. This is effective in highly challenging scenarios, e.g., maintaining the arrangement order of items on supermarket shelves. In such detail-oriented situations, adding cross-attention provides a rational consistency boost, at the cost of additional computational costs.

### v3

@@ 1 Introduction
Storytelling capability, narrative coherence, and context quality are increasingly important in the video generation community. The latest commercial models like Sora2 [46], Veo3.1 [20], and Kling2.5 [37] enable storyboard creation-based, scene planning, and dynamic camera shots (e.g., in the style of TikTok or YouTube Shorts). Recent academic models have also focused on long video streaming [72, 27], scene planning [22, 26, 83, 86, 87], and video context consistency [28, 31, 42]. Among these approaches, autoregressive models are a key paradigm for video storytelling, supporting native video continuation and storyboard streaming. Autoregressive models view video history as context, and they face unique challenges when handling long-form content like in-the-wild videos and movies.
The trade-off between context quality and length is a typical and critical problem in video context modeling. A naive sliding window, which cuts off all distant-enough frames, maintains a fixed context length but loses long-range history. Various compression approaches have been extensively studied; e.g., highly compressed VAEs like LTXV [23] and DC-AE [10] show that it is possible to compress pixel-space videos into more compact contexts, and hybrid approaches [1, 32, 88] like FramePack [78] show that compression can happen at multiple levels, albeit at the cost of high-frequency image details. Another strategy is to preserve the context length but reduce computation (e.g., using sparse [64, 77, 80, 65] or linear [5, 68, 58, 11, 74, 33] attention), though the linear layer cost remains an overhead for training and bidirectional inference. More specialized methods like token merging [4, 3] also demonstrate that a higher merging rate leads to greater detail loss. These research efforts implicitly converge on the observation that context quality and length are governed by a fundamental trade-off where the optimal balance has not yet been reached.
@@
Experiments show that the pretrained memory encoder can, by itself, reconstruct history frames at high quality (measured by PSNR/SSIM [60]). When connected to video diffusion models for autoregressive finetuning, the encoder provides long-range consistency. This framework achieves a balanced and optimized point in the context length-quality trade-off, leading to a practical autoregressive video model with a relatively short context length and perceptually high consistency based on its long-range history.
@@ 2 Related Work
Autoregressive video diffusion and long videos. Diffusion has become the driving force behind video synthesis, where a central challenge is length scaling. Training-free length‑extension methods [48, 43, 44, 84, 49] reschedule noise or re-balance temporal frequency to stretch pre-trained models beyond their training horizon. A complementary thread blends diffusion with causal prediction: Diffusion Forcing [7] and HistoryGuidance [52] enable variable-horizon conditioning and stable long rollouts by noise injection. These approaches are adapted in industrial systems such as SkyReels‑V2 [8] and Magi-1 [55]. Additionally, StreamingT2V [24] augments existing models with short- and long-term memory with randomized blending to generate longer videos. FAR [21] studies long‑context AR with flexible RoPE [53] decay and mixed short/long windows, while CausVid [72] distills a bidirectional teacher to a few‑step causal generator. StreamDiT [35] combines multi-step distillation with a moving frame buffer and mixed partition training to generate results in real-time. To mitigate error accumulation during AR generation, Self‑Forcing [27] simulates AR rollout during training, while its extensions [12, 41, 69] further improve length generalization.
@@
Context learning and compression. Long-context persistence is crucial as we extend video generation beyond a few seconds. One option is retrieval, which grounds generation in a persistent state. WorldMem [67] and Context-as-Memory [73] augment AR video world models with FoV-based history retrieval, while VMem [39] indexes past views by surfels to retrieve the most relevant viewpoints. Memory Forcing [25] pairs geometry-indexed spatial memory with tailored training regimes to balance exploration and revisits. Beyond fixed retrieval rules, Mixture-of-Contexts [6] learns a dynamic sparse attention context router so that tokens attend only to the most salient chunks. Similarly, MoGA [30] and Holocine [45] also propose token-level sparse attending policies. Pack-and-Force [63] instead proposes a learnable context semantic retriever. Orthogonal to retrieval, compression of history turns an unbounded context into a compact state. FramePack [78] compresses prior frames into a fixed‑size latent “packed” context. Captain Cinema [66] uses a similar compression for keyframes. StateSpaceDiffuser [50] and Po et al. [47] swap quadratic attention for recurrent states to maintain long‑term memory. TTTVideo [13] and LaCT [81] use light MLP test-time training layers as a learned context representation.
@@
Efficient video diffusion designs. As we scale video generation to long horizons, large context windows are bottlenecked, driving a wave of efficient computational designs. Kernel advances such as FlashAttention [14, 15] improve throughput. Static or hardware-friendly sparse patterns include sliding/tiling 3D windows [80], radial spatiotemporal masks [40], and training-free head/pruning heuristics [64, 70]. Dynamic or learned pruning/routing further select salient token pairs or blocks [62], coarse-to-fine sparse token selection [79], Sage/SpargeAttention families [76, 75, 77], blockified routing with cached search [65], and progressive block carving [82]. Another line of work leverages compressing the latent space or token sequence: token merging and patch scaling [2, 38], compact/variable-rate tokenizers [1], highly compressed latent space [23], or multiscale pyramids with re-noising [32]. SANA [68] introduced a linear‑attention Diffusion Transformer for images, while SANA‑Video [9] extends this with block‑linear attention and a constant‑memory KV cache.
@@ 3 Method
Preliminaries. We clarify the notation of a typical autoregressive video diffusion model. Unless otherwise noted, all “frames”, “pixels”, etc., refer to latent concepts. We consider the typical Diffusion Transformers (DiTs) like Wan [56] and HunyuanVideo [36] with flow match scheduling. Rectified-flow models map noisy latents $\bm{X}_{t_{i}}\in\mathbb{R}^{T\times H\times W\times C}$ with
@@
where $\bm{c}$ is a set of conditions such as text prompts (a single prompt or streamed from a storyboard), and $t_{i}\sim\mathcal{L}(0,1)$ is the shifted logit-normal distribution [18] for flow matching.
@@ 3.2 Finetuning Video Diffusion Model
Figure 5: Visual comparison on compression reconstruction. We present reconstruction results after the pretraining using different possible neural structures and various compression settings. * The “Large Patchifier” is technically equivalent to [78].
@@ 4.1 Experimental Details
Base models. All experiments are conducted on base models of HunyuanVideo [36] and Wan [56] family. For Wan 2.2, the involved models are the 5B model and the 14B high noise model. Only the high noise model is used to simplify experiments (similar to the design choice of WanAnimate).
Data preparation. The dataset is prepared with about 5 million internet videos from a range of websites. About half of the data is vertical Short-style videos and the rest are common horizontal videos. The data is cleaned for quality, and then the high-quality part is captioned with Gemini-2.5-flash VLM and the remainder is processed with local VLMs like QwenVL [57]. The generated captions are in storyboard format with timestamps. When training the autoregressive models, only the prompt with the nearest timestamp is received by the model.
Storyboards. If not noticed, all storyboard prompts in this paper are written by Gemini-2.5-pro. We also tested open source alternatives like Qwen [57] and the quality difference is negligible. See also the supplement for instructions.
@@ 4.2 Ablation Study
Table 1: Quantitative results on compression structures. We show numerical tests with different ablative compression architectures. * The “Large Patchifier” is technically equivalent to [78].
@@
Compression model architecture. As shown in Fig. 5, we compare the image reconstruction details of different architectures under various compression rates, focusing on the performance of Only LR, Without LR, Large Patchifier, and our proposed method. The Only LR variant keeps only the low-resolution branch in Fig. 3, while Without LR removes it, leaving only the high-resolution branch. The Large Patchifier method increases the compression rate by enlarging the DiT’s patchifying projection kernel without introducing new layers, equivalent to FramePack [78] and we directly use its implementation. The results indicate that the Only LR and Without LR variants produce reconstructions that differ significantly from the original images. The Large Patchifier method yields substantial structural changes. The proposed method, even at a high compression rate such as $4\times 4\times 2$, can still preserve much of the image appearance. A lower compression rate, such as $2\times 2\times 1$, provides the best detail preservation, albeit at the cost of a longer context length.
@@
Error accumulation (drifting). The problem of error accumulation highly depends on the training dataset. Surprisingly, we find that when trained on videos with dense camera shot shifts, like Short-style videos, the error accumulation problem seems to be absent or invisible. For other types of videos requiring long continuation over single shots, multiple design choices [78, 27, 7] are discussed in the supplement.
@@ 4.4 Quantitative Results
Metrics. We include several metrics for video evaluations from VBench [29], VBench2 [85], etc., with some modifications. Cloth: The cloth consistency by asking VLM questions. We upgrade VBench2 [85]’s LLaVA to Gemini-2.5-pro; Instance: The object-level quality by asking VLM questions about unnatural or inconsistent instances. We upgrade VBench2 [85]’s Qwen to Gemini-2.5-pro; Identity: The facial similarity using ArcFace [16] with face detection RetinaFace [17]; Clarity: The MUSIQ [34] image quality predictor trained on SPAQ [19]. This metric evaluates artifacts like noises; Aesthetic: The LAION aesthetic predictor [51], measuring the aesthetic values perceived by a CLIP estimator; Dynamic: The RAFT [54] modified by VBench to estimate the degree of dynamics. Note that multi-shot videos intuitively have a special RAFT flow between shots, but in practice, RAFT seems to output near-zero values between shots, making the difference negligible from regular videos; Semantic: The video-text score computed by ViCLIP [59]. This metric measures the overall semantic consistency and prompt alignments.
@@
WanI2V [56] + QwenEdit [61] (1p) 94.10 68.45 85.21 /
WanI2V [56] + QwenEdit [61] (2p) 95.09 68.22 91.19 1198
WanI2V [56] + QwenEdit [61] (3p) 94.28 67.30 80.34 /
@@
Consistency analysis. We compare various architectures of our method, including different compression rates and ablations with or without pre-training and the LR branch. We also compare against a common baseline that combines Wan-2.2-I2V [56] with Qwen-Image-Edit-2509 [61]. This baseline first generates a series of initial frames with an image model, animates them into video clips, and then concatenates the clips. We test this baseline with 1, 2, and 3 images for Qwen-Edit as history contexts (since the model supports at most 3 inputs). Our evaluation focuses on the consistency of clothing, identity, and objects. For several practical and well-performed models, we also conducted a user study and evaluated their ELO scores.
@@ 4.5 Alternative Architecture
HunyuanVideo [36] 12.8B ($\times4$$\times$2) 61.27 67.49 71.22 26.29 1189
Wan [56] 2.2 14B* ($\times4$$\times$2) 67.22 69.37 69.81 27.12 1231
Wan [56] 2.2 5B ($\times4$$\times$2) 66.25 69.01 65.13 25.99 1215
Wan [56] 2.2 5B ($\times2$$\times$2) 66.37 68.95 66.29 26.13 1224
@@
Cross attention enhancement. As shown in Fig. 9, we demonstrate that connecting the features from the penultimate layer of the encoder to each block of the DiT via cross-attention (like IP-Adapter [71]) can further enhance the model’s consistency. This is effective in highly challenging scenarios, e.g., maintaining the arrangement order of items on supermarket shelves. In such detail-oriented situations, adding cross-attention provides a rational consistency boost, at the cost of additional computational costs.